OpenStudy (anonymous):

If x>0, y>0, a x b, d y c, prove that: (a-d) |x| + |y| (b-c)

OpenStudy (anonymous):

\[a\leq x \leq b\] \[d \leq y \leq c\] \[x>0,y>0\] and you want \[a-d<|x| + |y| \leq b-c\] is that he problem?

OpenStudy (anonymous):

i'm sorry it's y < 0

OpenStudy (anonymous):

so maybe it would help if you wrote \[a-d<x-y<b-c\]? not sure but it could

OpenStudy (anonymous):

oh yes, you get it right away.

OpenStudy (anonymous):

just subtract. you have \[a<x<b\] \[d<y<c\] subtract gives \[a-d<x-y<b-c\] in one step.

OpenStudy (anonymous):

okay. and then?